HOw do you solve (y-4)^2=18
Answer by edjones(8007)

(y-4)^2=18
y-4=+-sqrt(18) take sqrt of each side.
y=4+-3sqrt(2)
y=4+3sqrt(2), y=4-3sqrt(2)
.
Ed